OpenStudy (anonymous):

what is derivative of 2x^3/2

OpenStudy (anonymous):

Recall the power rule: $$\dfrac{d}{dx}x^n = nx^{n-1}$$

OpenStudy (anonymous):

As well as the coefficient rule:$$\dfrac{d}{dx}C f(x) = C \dfrac{d}{dx} f(x) \text{ where }C \in \mathbb R$$

OpenStudy (anonymous):

is it 4/3 x^1/2

OpenStudy (anonymous):

I would check the coefficient on your derivative. Combining the two rules I gave above: $$\dfrac{d}{dx}Cx^n = Cnx^{n-1}$$

OpenStudy (anonymous):

Your exponent was correct, though.